Find the equation of the line through (4,3) perpendicular to the line 3x + y = 7
Answer by chibisan(131)

3x + y = 7
y = 7-3x

grad = negative reciprocal of -3 = 1/3

new equation :
y = 1/3x + b
sub(4,3)
3 = 1/3(4) + b
3 - 4/3 = b
5/3 = b

new equation will be :
y = 1/3x + 5/3
3y = x + 5(ans)